Assume in the market with two consumers, Chloe and Oliver. Chloe’s utility function is U(x, y) = x + 80y1/2. Oliver’s utility function is U(x, y) = x + 8y. At a Pareto optimal allocation in which both individuals consume some of each good, how much y does Chloe consume?

a. 25

b. 18

c. 121

d. We can’t tell without knowing the initial endowments.
Answer #1

a. 25

(At the pareto optimal allocation, MRS for both consumers is equal.
Chloe: MRS = MUx/MUy = (∂U/∂x)/(∂U/∂y) = 1/80(1/2)y(1/2)-1 = 1/40y-1/2 = y1/2/40
Oliver: MRS = MUx/MUy = (∂U/∂x)/(∂U/∂y) = 1/8

So, y1/2/40 = 1/8
So, y1/2 = 40/8 = 5
So, y = 52 = 25)